# Program Semantic Inequivalence Game with Large Language Models

### v1

@@ Abstract
In this work, we explore a method to synthetically generate code reasoning training data based on a semantic inequivalence game (SInQ): a generator agent creates program variants that are semantically distinct, derived from a dataset of real-world programming tasks, while an evaluator agent has to identify input examples that cause the original programs and the generated variants to diverge in their behaviour, with the agents training each other semi-adversarially. We prove that this setup enables theoretically unlimited improvement through self-play in the limit of infinite computational resources.
@@ 1 Introduction
Assistants based on Large Language Models (LLMs) are widely used by programmers to assist with coding tasks. While they perform well on common tasks, they still struggle with non-trivial reasoning about program semantics (Miceli Barone et al., 2023; Maveli et al., 2025). This limitation can lead to subtle bugs or prevent the detection of preexisting vulnerabilities and adversarial backdoors (Dinh et al., 2023; Dou et al., 2024), ultimately compromising the safety and security of generated code (Wang et al., 2024; Mohsin et al., 2024).
@@
Self-play involves training AI agents by pitting them against each other in adversarial games, incentivizing them to discover and defend against unusual scenarios. This approach has enabled AI systems to achieve human-level or even superhuman performance in games such as Go (Silver et al., 2016; 2017b), Chess (Silver et al., 2017a), Dota 2 (OpenAI et al., 2019), StarCraft II (Arulkumaran et al., 2019), and even social games involving dialogue, such as Diplomacy (FAIR, 2022). However, self-play methods typically require external engines to enforce game rules and compute scores, making them challenging to apply to open-ended tasks like coding. Recreational competitive coding environments such as CROBOTS [^2] https://github.com/tpoindex/crobots are overly domain-specific and impose strict performance limits, making them unsuitable for training agents in general code reasoning.
In this work, we introduce a game based on program semantic equivalence (or inequivalence) designed to train agents in code reasoning across arbitrary domains. By design, this game has no theoretical performance cap. We use it to train LLMs through self-play, demonstrating significant performance improvements on challenging tasks.
@@ 2 Proposed method
Our approach is based on the semantic equivalence of programs. This allows for precise verification of solutions, unlike problems based on natural language specifications or unit tests, which offer only partial coverage. Moreover, it is fundamentally linked to computability theory through reductions to Rice’s theorem and the Halting problem.
@@
2.1 The semantic equivalence game
Two programs $P$ and $Q$ are semantically equivalent if, for any given input $x$, they produce the same output $y$ and either both halt or neither does. Determining semantic equivalence is a fundamental problem in program verification and compiler design. It is undecidable in the general case as a direct consequence of Rice’s theorem (Rice, 1953), which implies that any deductive reasoning problem expressible in first-order logic can be reduced to it.
We define the semantic equivalence game as a one-shot interaction between two players: the generator, "Alice," and the evaluator, "Bob":
  1. 1.
    Alice receives a program $P$ and generates another program $Q$, which may be either equivalent or inequivalent to $P$.
  2. 2.
    If $P$ and $Q$ are equivalent, Alice must provide a machine-checkable proof. If they are inequivalent, Alice must supply a diverging input $x$ that causes them to produce different outputs or exhibit different halting behaviours.
  3. 3.
    The equivalence proof or diverging input is verified. If incorrect, Alice loses.
  4. 4.
    Bob receives $P$ and $Q$ and must determine whether they are equivalent. If Bob is correct, he wins and Alice loses; if he is wrong, Alice wins.
Alice’s objective is to generate challenging instances for Bob, while Bob’s goal is to solve them. Both agents improve iteratively through repeated play. The source programs $P$ provided to Alice are sampled from a dataset, such as a collection of short, self-contained programming exercises spanning a variety of tasks (e.g., MBPP (Austin et al., 2021)). This ensures the game remains grounded in real-world coding problems. If Alice were allowed to generate both $P$ and $Q$, she might develop an incentive to produce obscure, obfuscated code that does not contribute to Bob’s general reasoning skills.
Due to the undecidability of semantic equivalence, a perfect Bob cannot exist. This results in a "full-employment theorem" (Appel & Ginsburg, 1998) for Alice: in principle, she can always find new ways to fool Bob. This iterative process leads to increasingly stronger Bobs, who in turn train progressively stronger Alices. Unlike games such as Go or Chess, where perfect play is theoretically possible, the semantic equivalence game has no theoretical performance cap—agents can continue learning new coding logic indefinitely, constrained only by model capacity and computational resources.
The main drawback of the semantic equivalence game is its reliance on LLMs to generate formal equivalence proofs between programs. Most popular programming languages, such as Java and Python, are defined through natural language specifications or reference implementations rather than formal semantics. Even when formal semantics exist for a subset of a language, automatically generating fully machine-checkable equivalence proofs for non-trivial code is extremely challenging—likely beyond the capabilities of current LLMs and even most human programmers. We sidestep this issue by defining a different game that focuses solely on inequivalent programs.
@@ 2.2 The semantic inequivalence game
Similar to the equivalence game, we define the semantic inequivalence game as the following one-shot interaction between two players: the generator, "Alice," and the evaluator, "Bob":
@@
As before, Alice’s objective is to generate challenging instances for Bob, while Bob’s goal is to solve them. In this game, correctness can be verified simply by executing the programs on the provided diverging inputs, eliminating the need for generating and verifying formal proofs.
@@
To approximate non-termination detection, we impose a randomized time limit that significantly exceeds the typical runtime of source programs. This prevents Alice from exploiting a fixed time limit—for example, by generating a program $Q$ that loops for a predetermined duration before returning the same output as $P$.
@@
Although this game is more restricted than the semantic equivalence game, it has essentially the same logical complexity: given an infinite time limit, Bob’s task remains undecidable (see Appendix A). Thus, this game also has no theoretical performance cap, allowing agents to learn arbitrarily complex coding logic in principle while remaining grounded in a dataset of practical, real-world programming problems.
@@
The semantic inequivalence game is entirely adversarial and essentially a zero-sum game, provided that Alice generates only valid outputs. In some cases, modifying the game to be positive-sum may be beneficial—both to facilitate integration with supervised fine-tuning (SFT) and to prevent degenerate strategies (e.g., Alice generating excessively difficult cryptographic puzzles). We discuss these considerations further in Section 2.4 and Appendix B.
2.3 Implementation with Reinforcement Learning
The semantic inequivalence game, as defined above, is well-suited for reinforcement learning.
In a simple implementation, we prompt the Alice LLM with a program $P$ and generate a response that includes a chain-of-thought reasoning trace, a generated program $Q$, and a diverging input $x$. Alice receives a negative reward $r_{\text{Alice, invalid}}$ if $Q$ or $x$ cannot be correctly extracted from the output, or if $x$ is not a valid diverging example.
If Alice’s output is valid, we then prompt the Bob LLM with $P$ and $Q$ to generate an assessment response, including a chain-of-thought reasoning trace and a diverging input $x^{\prime}$. If Bob produces a well-formed response and $x^{\prime}$ is a valid diverging input, he receives a positive reward $r_{\text{Bob, good}}$, while Alice is penalized with $r_{\text{Alice, too easy}}$. If Bob makes a mistake, he is penalized with $r_{\text{Bob, bad}}$, while Alice is rewarded with $r_{\text{Alice, good}}$. These reward values are hyperparameters. [^3] Distinct rewards could also be assigned for syntactic errors and semantic mistakes.
While this approach is theoretically viable, it has several practical drawbacks:
  • The reward signal for Alice is relatively sparse, especially in early training stages, as Alice initially lacks information about Bob’s strength and thus generates easier instances.
  • Reinforcement learning is generally more expensive than supervised fine-tuning. It requires extensive hyperparameter tuning and is often unavailable to end users of API-only models, such as those from OpenAI.
Therefore, we instead adopt an implementation that relies on denser training signals and is more suitable for supervised fine-tuning.
@@ 2.4 Implementation with Supervised Fine-tuning and Difficulty Targeting
The difficulty value could be used directly as a more informative RL reward signal for Alice, but it also enables the following batched, supervised fine-tuning (SFT) implementation:
@@
We can continue to extend this dataset across multiple generations of Alice, as long as Bob remains unchanged. Once we update Bob (e.g., using SFT on its own successful input-output pairs), we need to recompute all the difficulty estimates for the programs in Alice’s dataset, as Bob is presumably stronger. We have found it beneficial to train Alice for many epochs before training a new Bob. Initially, Alice tends to generate examples that are too easy for Bob (especially when both Alice and Bob are based on the same LLM). Ideally, we would continue training Alice until convergence before each new Bob training run.
@@ 3.1 Training
[…]
@@ 3.3.2 Vulnerability Detection
Security vulnerabilities in code often arise from counterintuitive behaviours, where the intuitive understanding that programmers—whether human or LLM—have of the code’s semantics differs from its actual semantics in edge cases that evade pre-deployment testing. Our semantic inequivalence game incentivises the generator Alice to find edge cases that fool the evaluator Bob, who is then incentivised to become more robust by improving his reasoning about code semantics. Ideally, these capabilities should generalize to security vulnerability detection.
@@ 5 Limitations and Future Work
We release all the code necessary to reproduce our experiments, along with the synthetic training data we generated. Exact replication, limited by sampling randomness, should be possible with a modest budget (approximately \$250), as long as gpt-4o-mini-2024-07-18 remains available on the OpenAI platform.
@@
Antonio Valerio Miceli-Barone devised the initial research idea, performed the experiments and proved the theorems and drafted the initial version of this manuscript. Vaishak Belle and Ali Payani, advised the generation of the research idea, the execution of the experiments and helped revise this manuscript.
@@ Appendix A Non-decidability of semantic inequivalence
In the semantic equivalence game (section 2.1), the task of the evaluator ("Bob") of determining whether two programs $P$ and $Q$ are equivalent is clearly undecidable due to a trivial consequence of Rice’s theorem. But in the semantic inequivalence game (section 2.2), the programs $P$ and $Q$ given to Bob are guaranteed to be not equivalent, and Bob’s task is to find a diverging input $\hat{x}$ such that $P(\hat{x})\neq Q(\hat{x})$, which is guaranteed to exist. It may be asked whether this constraint on the programs makes the problem any logically easier. We show here that this is not the case.
@@
The mapping between programs and functions is surjective but not injective: each function can be defined by infinitely many programs. We define two programs $P$ and $Q$ equivalent if they define the same function, conversely we define them inequivalent if they define different functions, that is, if there exist at least one diverging input $\hat{x}$ such that $P(\hat{x})\neq Q(\hat{x})$.
@@
otherwise if $A$ does not halt on $n$, then $P_{A,n}^{*}$ and $Q_{A,n}^{*}$ diverge only on input 00:
@@ Appendix C Prompt templates
The evaluation prompts are included in the released code.

### v2

@@ Abstract
In this work, we explore a method to synthetically generate code reasoning training data based on a semantic inequivalence game (SInQ): a generator agent creates program variants that are semantically distinct, derived from a dataset of real-world programming tasks, while an evaluator agent has to identify input examples for which they behave differently. The agents train each other semi-adversarially, improving their ability to understand the underlying logic of code.
@@ 1 Introduction
Figure 1: Semantic inequivalence game: Alice receives a source program $P$ and generates a variant program $Q$ and a diverging input. Bob receives $P$ and $Q$ and generates another diverging input.
Assistants based on Large Language Models (LLMs) are widely used by programmers for coding tasks. While they perform well on common tasks, they still struggle with non-trivial reasoning about program semantics (Miceli Barone et al., 2023; Maveli et al., 2025). This limitation can lead to subtle bugs or prevent the detection of preexisting vulnerabilities and adversarial backdoors (Dinh et al., 2023; Dou et al., 2024), ultimately compromising the safety and security of generated code (Wang et al., 2024; Mohsin et al., 2024).
@@
Self-play involves training AI agents by pitting them against each other in adversarial games, incentivizing them to discover and defend against unusual scenarios. This approach has enabled AI systems to achieve human-level or even superhuman performance in games such as Go (Silver et al., 2016; 2017b), Chess (Silver et al., 2017a), Dota 2 (OpenAI et al., 2019), StarCraft II (Arulkumaran et al., 2019), and even social games involving dialogue, such as Diplomacy (FAIR, 2022). However, self-play methods typically require external engines to enforce game rules and compute scores, making them challenging to apply to open-ended tasks like coding. Recreational competitive coding environments such as CROBOTS [^2] https://github.com/tpoindex/crobots are overly domain-specific and impose strict performance limits, making them unsuitable for training agents in general code reasoning. We are aware of only one work, Zhao et al. (2025), concurrent to our own, which uses self-play to train LLMs for arbitrary code generation, while Dong & Ma (2025) use a similar approach for theorem proving.
In this work, we introduce a game based on program semantic inequivalence designed to train agents in code reasoning across arbitrary domains. By design, this game has no theoretical performance cap. We use it to train LLMs through self-play, demonstrating significant performance improvements on challenging tasks.
@@ 2 Proposed method
Our approach is based on the semantic equivalence of programs, or more specifically, semantic inequivalence. This allows for precise verification of solutions, unlike problems based on natural language specifications or unit tests, which offer only partial coverage. Moreover, it is fundamentally linked to computability theory through reductions to Rice’s theorem and the Halting problem.
@@ 2.1 The semantic inequivalence game
Two programs $P$ and $Q$ are semantically equivalent if, for any given input $x$, either they both halt producing the same output $y$ or neither halts. Determining semantic equivalence is a fundamental problem in program verification and compiler design, but automatic proving equivalence between arbitrary programs is complicated since popular programming languages, such as Java and Python, are defined through natural language specifications or reference implementations rather than formal semantics. Even when formal semantics exist for a subset of a language, automatically generating fully machine-checkable equivalence proofs for non-trivial code is quite challenging even for expert human programmers. We sidestep this issue by defining a program understanding game that focuses solely on inequivalent programs.
We define the Semantic Inequivalence Game as the following one-shot interaction between two players: the generator, "Alice," and the evaluator, "Bob":
@@
Alice’s objective is to generate challenging instances for Bob, while Bob’s goal is to solve them. In this game, correctness can be verified simply by executing the programs on the provided diverging inputs, eliminating the need for generating and verifying formal proofs.
Both agents are trained iteratively through repeated play. The source programs $P$ provided to Alice are sampled from a dataset, such as a collection of short, self-contained programming exercises spanning a variety of tasks (e.g., MBPP (Austin et al., 2021)). This ensures that the game stays grounded to real-world coding problems. If Alice were allowed to generate both $P$ and $Q$, she might develop an incentive to produce unusual, obfuscated code that might not contribute to Bob’s general reasoning skills.
@@
To approximate non-termination detection, we impose a randomized time limit that significantly exceeds the typical runtime of source programs. This prevents Alice from exploiting a fixed time limit, for example, by generating a program $Q$ that loops for a predetermined duration before returning the same output as PP. [^3] The time limit is randomized to discourage Alice from gaming the system by crafting artificial delays, which could lead to uninteresting cases.
Example 1.
1. Alice receives program $P$:
def fib(n):
if n <= 0:
return 0
elif n == 1:
return 1
return fib(n - 1) + fib(n - 2)
and returns program $Q$:
def fib(n):
if n == 0:
return 0
elif n == 1:
return 1
return fib(n - 1) + fib(n - 2)
together with diverging input $x$:
{"n": -1}
2. Both programs are run in a sandbox with input $x$, which results in $P$ returning 0 while $Q$ recurs until it triggers either the recursion limit of the Python interpreter or the randomized time limit, proving that $x$ is indeed a diverging input.
3. Bob receives both $P$ and $Q$ a generates a possibly different diverging input $\hat{x}$, e.g.:
{"n": -2}
$P$ and $Q$ are evaluated on input $\hat{x}$, proving that this is also a diverging input, therefore, Bob wins this round.
@@ 1. Alice receives program $P$:
Unlike games such as Go or Chess, where perfect play is theoretically possible, the semantic inequivalence game has no strict performance cap: given an infinite time limit, Bob’s task is undecidable (see Appendix A). This implies that in principle the agents can learn arbitrarily complex coding logic while remaining grounded in a dataset of practical, real-world programming problems.
@@
The semantic inequivalence game is entirely adversarial and essentially a zero-sum game, provided that Alice generates only valid outputs. In some cases, modifying the game to be positive-sum may be beneficial, both to facilitate integration with supervised fine-tuning (SFT) and to prevent degenerate strategies (e.g., Alice generating excessively difficult cryptographic puzzles). We discuss these considerations further in Appendix B.
@@ 2.2 Implementation with Supervised Fine-tuning and Difficulty Targeting
The semantic inequivalence game, as defined above, is well-suited for reinforcement learning, however, reinforcement learning was not available on the OpenAI API at the time of our experiments, therefore, we devised the following rejection sampling fine-tuning implementation, with explicit difficulty supervision.
@@
We can continue to extend this dataset across multiple generations of Alice, as long as Bob remains unchanged. Once we update Bob (using rejection sampling SFT on its own successful input-output pairs), we need to recompute all the difficulty estimates for the programs in Alice’s dataset, as Bob is presumably stronger. We have found it beneficial to train Alice for many epochs before training a new Bob. Initially, Alice tends to generate examples that are too easy for Bob (especially when both Alice and Bob are based on the same LLM). Ideally, we would continue training Alice until convergence before each new Bob training run.
@@ 3.1 Training
Both Alice and Bob are instructed to produce output in markdown format, using markdown sections to distinguish their CoT traces from the final outputs, which are embedded in Python code blocks. We sample from the models with a temperature of $1.0$ and $\text{top\_p}=0.7$, generating $N=10$ samples per query. We use the Mistune markdown parser [^4] https://mistune.lepture.com/en/latest/ followed by the Python ast parser/unparser. This step extracts, syntactically validates, and normalizes the outputs [^5] Parsing and then unparsing the Python code with ast removes comments or unusual indentation styles, preventing trivial non-semantic attacks.. We then semantically check the diverging inputs against the pairs of source and generated programs by executing them within a test harness, using a randomized time limit, uniformly sampled between 2.5 and 5.5 seconds, to discourage Alice from generating instances that rely on race conditions.
@@ 3.3.2 Vulnerability Detection
Security vulnerabilities in code often arise from counterintuitive behaviours, where the intuitive understanding that programmers, whether human or LLM, have of the code’s semantics differs from its actual semantics in edge cases that evade pre-deployment testing. Our semantic inequivalence game incentivises the generator Alice to find edge cases that fool the evaluator Bob, who is then incentivised to become more robust by improving his reasoning about code semantics. Ideally, these capabilities should generalize to security vulnerability detection.
@@ 5 Limitations and Future Work
We will release all the code necessary to reproduce our experiments, along with the synthetic training data we generated, upon publication. Exact replication, limited by sampling randomness, should be possible with a modest budget (approximately \$250), as long as gpt-4o-mini-2024-07-18 remains available on the OpenAI platform.
@@
Vaishak Belle and Ali Payani, advised the generation of the research idea, the execution of the experiments and helped revise this manuscript.
@@ Acknowledgments
Additional funding was provided by the UKRI Engineering and Physical Sciences Research Council and the Artificial Intelligence Security Institute under the grant "Understanding and Improving the Behaviour of AI Agents in Competitive and Cooperative Games".
@@ Appendix A Non-decidability of semantic inequivalence
In a semantic equivalence game where the task of the evaluator ("Bob") is to determine whether two programs $P$ and $Q$ are equivalent, there is clear undecidablity due to a trivial consequence of Rice’s theorem (Rice, 1953). Since a perfect Bob cannot exist, this results in a "full-employment theorem" (Appel & Ginsburg, 1998) for Alice: in principle, she can always find new ways to fool Bob. This iterative process leads to increasingly stronger Bobs, who in turn train progressively stronger Alices.
[…]
@@
[…]
@@
otherwise if $A$ does not halt on $n$, then $P_{A,n}^{*}$ and $Q_{A,n}^{*}$ diverge only on input 0:
@@ Appendix C Prompt templates
The evaluation prompts will be included in the code released upon publication.